OpenStudy (anonymous):

How do you solve -3 over 2 squared?? (fraction)

OpenStudy (anonymous):

do 2^2=4 -3 over 4

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

I dont understand...... its -3/4 to the power of 2 so what do you do when you square fractions?

OpenStudy (anonymous):

What is the original problem?

OpenStudy (anonymous):

\[-3/2^2\]

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

2^2=4

OpenStudy (anonymous):

so...the answer is -3/4

OpenStudy (anonymous):

you do not apply the exponent to the numerator.

OpenStudy (anonymous):

So (-3/2)to the second power is -3/4? thx

OpenStudy (anonymous):

\[(-3/2)^2\]

OpenStudy (anonymous):

that is different than \[(-3/2^2)\]

OpenStudy (anonymous):

which one is it? the first or second one?

OpenStudy (anonymous):

1st one

OpenStudy (anonymous):

for that one, you take 3 to the second power, which is 9, then you take 2 to the second power, which is 4. So that anwser would be -9/4

OpenStudy (anonymous):

wouldn't it be 9/4 and not -9/4 because its -3squared

OpenStudy (anonymous):

yes :) sorry

OpenStudy (anonymous):

thx.... ur smart

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

i have another question... how do you factor r squared minus 3 over 2r + 9 over 4